Scores on a certain standardized test have a mean of 500, and a standard deviation of 100. How common is a score between 600 and 700? Calculate the probability.


1
Expert's answer
2021-10-27T14:50:50-0400

"\mu=500 \\

\sigma= 100 \\

P(600<X<700) = P(X<700) -P(X<600) \\

= P(Z< \frac{700-500}{100}) -P(Z< \frac{600-500}{100}) \\

= P(Z< 2) -P(Z< 1) \\

= 0.9772 -0.8413 \\

= 0.1359"